# PersonaFlow: Boosting Research Ideation with LLM-Simulated Expert Personas

### v1

@@ Abstract.
Developing novel interdisciplinary research ideas often requires discussions and feedback from experts across different domains. However, obtaining timely inputs is challenging due to the scarce availability of domain experts. Recent advances in Large Language Model (LLM) research have suggested the feasibility of utilizing LLM-simulated expert personas to support research ideation. In this study, we introduce PersonaFlow, an LLM-based system using persona simulation to support the ideation stage of interdisciplinary scientific discovery. Our findings indicate that using multiple personas during ideation significantly enhances user-perceived quality of outcomes (e.g., relevance of critiques, creativity of research questions) without increasing cognitive load. We also found that users’ persona customization interactions significantly improved their sense of control and recall of generated ideas. Based on the findings, we discuss highlighting ethical concerns, including potential over-reliance and cognitive biases, and suggest design implications for leveraging LLM-simulated expert personas to support research ideation when human expertise is inaccessible.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Natural language processing
@@ 1. Introduction
The development of novel research ideas often requires discussion and feedback collection with peer researchers (Al Qunayeer, 2020; Saeed et al., 2021; Man et al., 2018), especially for researchers conducting research on interdisciplinary topics. Researchers who are unfamiliar with a research domain often require guidance from domain experts to quickly gain an overview of the field. However, this requires timely and iterative input from experts, and can often be challenging due to the constraints of accessibility and availability of experts willing to provide feedback (Chugh et al., 2022). Meanwhile, feedback from peers with different backgrounds can be hard to interpret without domain knowledge, thus leading to communication gaps (August et al., 2024) between researchers from various disciplines when discussing research ideas. Moreover, researchers sometimes perceive expert critique as personal judgment, potentially leading to increased caution or defensiveness when seeking feedback (Wang and Li, 2011).
@@
Prior research has shown that AI can be used to deliver non-judgemental immediate feedback and interactions (Lee et al., 2020; Gratch et al., 2014; Ta et al., 2020). The concept of persona has been historically used in scenarios like design ideation (Chen et al., 2011; So and Joo, 2017; Salminen et al., 2020) to facilitate need-finding. The use of AI-based personas for feedback gathering from an audience perspective has been widely discussed in existing works under various ideation contexts such as design(Park et al., 2022; Li et al., 2024a), journalism (Petridis et al., 2023) and group-based decision-making (Chiang et al., 2024) as beneficial for ideation by simulating human thinking. Providing diverse perspectives using AI-simulated personas would analogically benefit interdisciplinary research ideation. However, scientific ideation (Foster, 2004) differs from design ideation, which often requires extensive domain knowledge and grounding in literature. It remains under-studied whether and how AI personas can be used as expert personas and provide knowledge-intensive feedback and opinions. In the context of using LLM for scientific ideation, prior studies have explored using LLMs to assist researchers in developing research ideas in domains such as HCI (Liu et al., 2024), ML/NLP (Wang et al., 2023a) and Economics (Korinek, 2023). While recent research has shown that LLMs have the capability to simulate personas with different expertise and effectively boost their performance in both knowledge- and reasoning-demanding tasks (Wang et al., 2023c; Chan et al., 2023; Li et al., 2023), there appears to be limited research specifically exploring the use of LLMs to simulate expert personas for facilitating research ideation in multidisciplinary fields. Thus, we aim to provide insights into the research question: how do researchers use AI-simulated expert personas to support their research ideation process when conducting interdisciplinary research?
@@
To this end, our study proposed a novel system, PersonaFlow, that aims to provide novel insights into how AI can help facilitate the peer feedback process through persona simulation. As shown in Figure 1, PersonaFlow is an LLM-based research ideation system that facilitates human-AI co-creation between human researchers and LLM-simulated expert personas. The system utilized a graph-based design that enables users to iteratively co-create and scaffold research ideas and questions. The system also adopts a mind map-styled to visually organize information and ideation outcomes. The design of the system is informed by a formative interview study with 10 researchers from three different research domains.
To evaluate PersonaFlow, we first conducted a task-based user study involving 21 researchers from 8 different research institutions to assess the effectiveness of the system. The study focused on evaluating users’ perceptions of and interactions with these personas. We also conducted an open-ended exploration study where 10 participants were invited to freely explore the system without tasks or constraints, to uncover their in-situ use cases of PersonaFlow in supporting the ideation process.
@@
The contribution of this study to the HCI community and the literature on human-AI co-creation is three-fold:
@@
  • First, this study introduced a novel LLM-based research ideation system designed to facilitate human-AI collaboration in generating research ideas through graph-based interactions. The system breaks down the interdisciplinary collaborative ideation process into four key steps, each represented by a type of node: the framing of research questions (RQ Node), seeking help from domain experts (Persona Node), gathering literature recommendations (Literature Node), and feedback (Critique Node) from domain experts. The system supports the automatic generation of follow-up nodes, such as suggesting Literature Nodes based on an RQ Node and a Persona Node, while also allowing users to fully customize the nodes and create their unique combinations. PersonaFlow addresses the challenge of limited access to domain experts in unfamiliar fields by leveraging LLM-simulated expert personas to facilitate interdisciplinary research ideation.
  • Second, our task-based evaluation provided empirical insights into users’ perceptions and behaviors. Specifically, users’ ability to edit personas enhanced their sense of control and facilitated their recall of ideas from unfamiliar domains. Additionally, the open-ended exploratory study demonstrated the system’s versatility and effectiveness in supporting users to either converge on a specific topic or explore diverse research questions. The expert personas also increased users’ trust by providing both relevant and unexpectedly insightful literature, encouraging deeper exploration.
  • Third, we offer novel insights by discussing our findings in relation to past research, particularly on how the integration of expert personas can not only improve ideation outcomes, but also enhance users’ knowledge acquisition and sense of agency. Additionally, we provide design implications for future research ideation systems that apply LLMs, highlighting the need of mitigating potential biases (e.g., customizing expert personas) during human-AI collaboration.
@@ 2. Related Work
In this section, to show how our work contributes to the existing literature, we provide a detailed review of relevant existing works from three perspectives. First, we discuss how LLM-based persona simulation has been employed for both task-oriented and creative applications. Then, we examine how LLMs have been used to support research ideation. Finally, we review prior literature on cognitive theories that inform the dynamics of collaborative ideation.
2.1. Persona simulation using LLM
@@
2.2. LLM for research ideation and scaffolding
Researchers have started exploring methods to use LLMs to augment research ideation and brainstorming processes for scientific writing (Baek et al., 2024; Wang et al., 2023a; Shen et al., 2023; Korinek, 2023). For instance, Nigam et al. (2024) employed an LLM-Agentic workflow that utilizes Retrieval-Augmented Generation (RAG) to identify similar papers based on an initial research topic provided by the user, utilizing colleague and mentor profile-based agents to perform motivation validation and method synthesis iteratively to develop a research proposal. Similarly, Gu and Krenn (2024) used a knowledge graph to identify sub-graphs relevant to a researcher’s interest and uses these sub-graphs alongside a model to suggest scientific research ideas based on initial research topics given by researchers. Guo et al. (2024) explored the usage of AI assistants on user ideation, specifically, analyzing the impact of AI to introduce bias in user ideation. Studies (Memmert and Tavanapour, 2023; Gonzalez et al., 2024; Shaer et al., 2024) have also discussed the use of LLMs for augmenting ideation and brainstorming processes for creative and group ideation. Other works, such as Sandholm et al. (2024)’s work employed a non-linear thinking methodology that utilizes a problem-solution dataset and fine-tuned LLMs to infer new problem statements from the original user problem statements. Lozano et al. (2023)’s work used RAG and fine-tuned LLMs to build a question-answering system using PubMed data for the medical domain. In this study, we explore the feasibility of using LLM-based techniques to support human-in-the-loop research ideation through the simulation of expert personas.
@@ 2.3. Cognitive theories for collaborative ideation
Scientific discovery requires both extensive knowledge and creative thinking processes (Langley, 1987; Aschauer et al., 2022; Foster, 2004). Research has sought to unveil the underlying cognitive mechanisms involved in scientific creativity, aiming to understand how scientists generate innovative ideas and make breakthroughs in their respective fields. Previous studies have highlighted the importance of various thinking processes, such as critical thinking (Newton, 2010), divergent thinking(Sak and Ayas, 2013; Hu et al., 2010), and convergent thinking (Agnoli et al., 2016), in fostering scientific creativity.
@@
The process of research ideation often also involves collaborative efforts, such as brainstorming sessions, peer discussions, and interdisciplinary collaboration, which can significantly enhance the potential for novel discoveries and innovative solutions (Siangliulue et al., 2015). On the other hand, studies (Ulrich and Eppinger, 2016; Kornish and Ulrich, 2011; Terwiesch and Ulrich, 2009) have examined the approach of parallel search in a group-based ideation setting, both in teams and as individuals, whose findings emphasize the importance of incorporating individual contributions that are independent of each other to improve ideation quality. Relevant to understanding the collaboration dynamics between members during group-based ideation, Transactive Memory System (TMS) theory (Lewis and Herndon, 2011) has been proposed as a theoretical framework for understanding the information dynamics in team-based collaboration, where the three main factors include specialization, coordination and credibility. Empirical studies (Ali et al., 2019; Dastmalchi et al., 2021; Hanke, 2006) have suggested the benefits of a well-developed transactive memory system for collaborative ideation in a team setting. Closely related, studies (Mayo et al., 2017; Harrison and Klein, 2007) also highlight the importance of diversity in team members’ expertise and background in promoting creative problem-solving and enhancing the overall efficacy of the group ideation process. Past research has sought to understand users’ behavior of interacting with AI personas under the context of writing (Benharrak et al., 2024) and chit-chatting (Ha et al., 2024). In this study, we also aim to deepen the understanding of users’ persona customization behavior for scientific ideation through empirical observations.
@@ 3.1. Formative Study: Informing the Design of PersonaFlow
All of the interviewees mentioned the challenges in interdisciplinary learning, including difficulties in understanding methods and jargon from other domains. Identifying and framing meaningful research problems in unfamiliar domains was also deemed challenging to participants, which involves selecting suitable methods and understanding their relevance and application in the new context (“… we always use like different kind of new research methods … all these kind of things were new to me, which is a big challenge for me … ,” P19). Participants also noted the challenge of finding and understanding relevant past research articles. The rapidly evolving nature of some fields (e.g., HCI and Biomedical), makes it challenging to find comprehensive and up-to-date literature. We also identified challenges in accessing domain-specific knowledge from experts. When seeking advice from peers, professors, or domain experts without already having an understanding of the domain, the guidance researchers receive is typically limited to being directed to research papers or resources, rather than receiving detailed explanations (“… it can be kind of hard to reach out and then try to explain what I want … some of the concepts require a lot of background knowledge and previous experience that I feel like they can’t really communicate to me just through an email, or like a 5 minutes talk … ,” P12). Participants also highlighted difficulties in finding the right person to consult, and there is often a gap in understanding between them and the experts they approach, as noted by P7 “… a gap of understanding between the two (researchers) … professors have different kind of understanding and expectation about our capabilities, they might be thinking that I might be aware of a statistical concept … but I do not know how to do it, and they’re like, but I thought you knew it already … .”
@@
  • [D2] Feedback gathering from the perspectives of domain experts: As highlighted by the formative study interviewees, in real-life research settings, getting feedback and sharing knowledge from domain experts is crucial for advancing novel ideas, but the process is often hindered by the lack of access and opportunities for communication to obtain timely and constructive input. Past theories (Mayo et al., 2017) have also suggested the significance of having access to the knowledge of a diverse set of expertise in the context of collaborative ideation. The system should have mechanisms for actively seeking and incorporating feedback from domain experts with diverse background knowledge in a timely manner.

### v2

# PersonaFlow: Designing LLM-Simulated Expert Perspectives for Enhanced Research Ideation

@@ Abstract.
Generating interdisciplinary research ideas requires diverse domain expertise, but access to timely feedback is often limited by the availability of experts. In this paper, we introduce PersonaFlow, a novel system designed to provide multiple perspectives by using LLMs to simulate domain-specific experts. Our user studies showed that the new design 1) increased the perceived relevance and creativity of ideated research directions, and 2) promoted users’ critical thinking activities (e.g., interpretation, analysis, evaluation, inference, and self-regulation), without increasing their perceived cognitive load. Moreover, users’ ability to customize expert profiles significantly improved their sense of agency, which can potentially mitigate their over-reliance on AI. This work contributes to the design of intelligent systems that augment creativity and collaboration, and provides design implications of using customizable AI-simulated personas in domains within and beyond research ideation.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; ; booktitle: DIS 25price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Interactive systems and toolsccs: Computing methodologies Natural language processing
@@
[…]
An overview of the PersonaFlow system’s interface design. Two panel diagram: Upper panel demonstrates a graph-based design with four types of nodes to mimic the process of gathering feedback on research ideas from AI-simulated expert personas. The design starts with users creating Research Question (RQ) Nodes, where users input their initial research topics. Persona Nodes represent AI personas that suggest relevant literature pulled from online databases, demonstrated as Literature Nodes. These Literature Nodes are connected to Critique Nodes which provide critiques and feedback. These critiques help users iteratively refine their initial RQs into revised RQ Nodes (5). Through this process, users can explore and improve their research questions until they are satisfied with the results. The lower panel of the diagram illustrates the different node categories and corresponding illustrations to signal the node function
@@ 1. Introduction
The development of novel research ideas often requires discussion and feedback collection with peer researchers (Al Qunayeer, 2020; Saeed et al., 2021; Man et al., 2018), especially for researchers conducting research on interdisciplinary topics. Researchers who are unfamiliar with a research domain often require guidance from domain experts to quickly gain an overview of the field. However, this requires timely and iterative input from experts, and can often be challenging due to the constraints of experts’ availability willing to provide feedback (Chugh et al., 2022). Meanwhile, feedback from peers with different backgrounds can be hard to interpret without domain knowledge, thus leading to communication gaps (August et al., 2024) between researchers from various disciplines when discussing research ideas. Moreover, researchers sometimes perceive expert critique as personal judgment, potentially leading to increased caution or defensiveness when seeking feedback (Wang and Li, 2011).
@@
Prior research has shown that AI can be used to deliver non-judgemental immediate feedback and interactions (Lee et al., 2020; Gratch et al., 2014; Ta et al., 2020). The concept of persona has been historically used in scenarios like design ideation (Chen et al., 2011; So and Joo, 2017; Salminen et al., 2020) to facilitate need-finding. The use of AI-based personas for feedback gathering from an audience perspective has been widely discussed in existing works under various ideation contexts such as design (Park et al., 2022; Li et al., 2024a), journalism (Petridis et al., 2023) and group-based decision-making (Chiang et al., 2024) as beneficial for ideation by simulating human thinking. Similarly, the use of AI-simulated personas may also benefit interdisciplinary research ideation when used for exploration of diverse perspectives or addressing blind spots due to the limitation of human expertise. However, scientific ideation (Foster, 2004) differs from design ideation, which often requires extensive domain knowledge and grounding in literature. It remains under-studied whether and how AI personas can be used as expert personas and provide knowledge-intensive feedback and opinions. In the context of using LLM for scientific ideation, prior studies have explored using LLMs to assist researchers in developing research ideas in domains such as HCI (Liu et al., 2024), ML/NLP (Wang et al., 2023a) and Economics (Korinek, 2023). While recent research has shown that LLMs have the capability to simulate personas with different expertise and effectively boost their performance in both knowledge- and reasoning-demanding tasks (Wang et al., 2023c; Chan et al., 2023; Li et al., 2023), there appears to be limited UX research exploring the use of LLMs to simulate the diverse perspectives of multidisciplinary experts with different backgrounds for facilitating research ideation, and how the different perspectives impact researchers’ ideation processes.
@@
To this end, our study proposed a novel system, PersonaFlow, that aims to provide novel insights into how AI can help facilitate the peer feedback process through simulating diverse domain experts. As shown in Figure 1, PersonaFlow is an LLM-based research ideation system that facilitates human-AI co-creation between human researchers and LLM-simulated experts. Informed by a formative interview study, the novel design utilized a graph-based design that enables users to iteratively co-create and scaffold research ideas and questions. The system also adopts a mind map-styled design to visually organize information and ideation outcomes. To evaluate PersonaFlow, we conducted a two-phase study. First, we ran a task-based study with 21 participants, evaluating users’ perceptions of and interactions with these experts, and examining how these changed temporally as they engaged with more expert perspectives temporally during the study. Second, we conducted an optional open-ended exploration session where 10 participants were invited to freely explore the system without tasks or constraints, to uncover their in-situ use cases of PersonaFlow in supporting the ideation process.
@@
The contribution of this paper to the HCI community and the literature on human-AI co-creation is three-fold:
@@
  • First, we designed, developed, and evaluated a novel LLM-based system, called PersonaFlow, which aims to facilitate human-AI collaboration in research ideation through graph-based interactions. The system supports interdisciplinary research ideation through four key steps, each associated with a specific node type: 1) framing research questions (RQ Node), 2) seeking help from domain experts (Persona Node), 3) gathering literature recommendations (Literature Node), and 4) obtaining feedback (Critique Node) from domain experts.
  • Second, our user evaluation studies provided new empirical evidence, demonstrating its effectiveness in supporting research ideation. Specifically, users perceived multiple perspectives offered by the simulated experts as more relevant and insightful. The design also enabled users to edit experts’ profiles, enhancing their sense of agency.
  • Third, the case studies illustrated the system’s versatility in supporting users’ critical thinking activities, e.g., interpretation, analysis, evaluation, inference, and self-regulation. For example, we observed that participants engaged in inference thought processes after receiving new relevant papers recommended by the (Literature Node); and they exhibited self-regulation by rephrasing their initial RQ after reviewing the (Critique Nodes) from different AI-simulated experts.
  • Fourth, we proposed design implications for mitigating potential confirmation and social biases that were revealed during participants’ use of the system. We further discussed the broader applications of such design platforms for scaffolding ideas from multiple perspectives in other domains, e.g., service design and creative writing.
@@ 2. Related Work
In this section, to show how our work contributes to the existing literature, we provide a detailed review of relevant existing works from three perspectives. First, we present how LLMs have been used to support research ideation. Then, we show the feasibility of LLM-based persona simulation to support both task-oriented and creative applications. Finally, we review prior literature on cognitive theories that inform the dynamics of collaborative ideation.
2.1. LLM for Research Ideation and Scaffolding
Researchers have started exploring methods to use LLMs to augment research ideation and brainstorming processes for scientific writing (Baek et al., 2024; Wang et al., 2023a; Shen et al., 2023; Korinek, 2023). For instance, Nigam et al. (2024) employed an LLM-Agentic workflow that utilizes Retrieval-Augmented Generation (RAG) to identify similar papers based on an initial research topic provided by the user, utilizing colleague and mentor profile-based agents to perform motivation validation and method synthesis iteratively to develop a research proposal. Similarly, Gu and Krenn (2024) used a knowledge graph to identify sub-graphs relevant to a researcher’s interest and uses these sub-graphs alongside a model to suggest scientific research ideas based on initial research topics given by researchers. Guo et al. (2024) explored the usage of AI assistants on user ideation, specifically, analyzing the impact of AI to introduce bias in user ideation. Studies (Memmert and Tavanapour, 2023; Gonzalez et al., 2024; Shaer et al., 2024) have also discussed the use of LLMs for augmenting ideation and brainstorming processes for creative and group ideation. Other works, such as Sandholm et al. (2024)’s work employed a non-linear thinking methodology that utilizes a problem-solution dataset and fine-tuned LLMs to infer new problem statements from the original user problem statements. Lozano et al. (2023)’s work used RAG and fine-tuned LLMs to build a question-answering system using PubMed data for the medical domain. In this study, we explore the feasibility of using LLM-based techniques to support human-in-the-loop research ideation through the simulation of expert personas. Prior systems aimed to automate research ideation or offer interaction through a single conversational assistant or agent, while few provided support to explore knowledge from multiple diverse perspectives. The PersonaFlow system fills this gap by enabling users to engage with diverse perspectives represented by LLM-simulated customizable expert personas. We show that through the user study the use of such expert perspectives can improve ideation quality and also stimulate critical‑thinking behaviors without added cognitive burden.
2.2. Simulating Diverse Perspectives using LLMs
@@ 2.3. Cognitive Theories for Collaborative Ideation
Scientific discovery requires both extensive knowledge and creative thinking processes (Langley, 1987; Aschauer et al., 2022; Foster, 2004). Research has sought to unveil the underlying cognitive mechanisms involved in scientific creativity, aiming to understand how scientists generate innovative ideas and make breakthroughs in their respective fields. Previous studies have highlighted the importance of various thinking processes in fostering scientific creativity, such as critical thinking (Newton, 2010), divergent thinking(Sak and Ayas, 2013; Hu et al., 2010), and convergent thinking (Agnoli et al., 2016). Moreover, critical thinking has been extensively examined within the domain of science education (Santos, 2017; Bailin, 2002; Vieira et al., 2011), establishing it as a crucial skill for development especially for novice researchers and those interested in engaging with new fields. Past research indicates that exposure to different perspectives and opinions improves critical thinking (Walton, 1989; Kuhn, 2019) during scientific research (Nussbaum et al., 2012; Hendriks et al., 2020). Critical thinking skills can also be fostered through the encouragement of perspective-taking (Southworth, 2022) and the integration of diverse viewpoints (Lai et al., 2011).
@@
The process of research ideation often also involves collaborative efforts, such as brainstorming sessions, peer discussions, and interdisciplinary collaboration, which can significantly enhance the potential for novel discoveries and innovative solutions (Siangliulue et al., 2015). On the other hand, studies (Ulrich and Eppinger, 2016; Kornish and Ulrich, 2011; Terwiesch and Ulrich, 2009) have examined the approach of parallel search in a group-based ideation setting, both in teams and as individuals, whose findings emphasize the importance of incorporating individual contributions that are independent of each other to improve ideation quality. Relevant to understanding the collaboration dynamics between members during group-based ideation, Transactive Memory System (TMS) theory (Lewis and Herndon, 2011) has been proposed as a theoretical framework for understanding the information dynamics in team-based collaboration, where the three main factors include specialization, coordination and credibility. Empirical studies (Ali et al., 2019; Dastmalchi et al., 2021; Hanke, 2006) have suggested the benefits of a well-developed transactive memory system for collaborative ideation in a team setting. Closely related, studies (Mayo et al., 2017; Harrison and Klein, 2007) also highlight the importance of diversity in team members’ expertise and background in promoting creative problem-solving and enhancing the overall efficacy of the group ideation process. Past research has sought to understand users’ behavior of interacting with AI-simulated experts under the context of writing (Benharrak et al., 2024) and chit-chatting (Ha et al., 2024). Building on the literature reviewed above, we conceptualize research ideation as a four‑phase cycle of divergent exploration, evidence gathering, feedback gathering, and convergent consolidation. We later designed PersonaFlow to support each of the four phases. In this study, we also aim to deepen the understanding of users’ expert profile customization behavior for scientific ideation through empirical observations. We examine the empirical results through the lens of TMS and critical thinking theories, thus providing implications of designs using LLM-simulated personas in the context of ideation. Our work hence extends prior LLM‑assisted ideation tools and explores a transferable design framework for future systems targeting different stages of ideation.
@@ 3.1. Formative Study: Informing the Design of PersonaFlow
All of the interviewees mentioned the challenges in interdisciplinary learning, including difficulties in understanding methods and jargon from other domains. Identifying and framing meaningful research problems in unfamiliar domains was also deemed challenging to participants, which involves selecting suitable methods and understanding their relevance and application in the new context (“… we always use like different kind of new research methods … all these kind of things were new to me, which is a big challenge for me … ,” P19). Participants also noted the challenge of finding and understanding relevant past research articles. The rapidly evolving nature of some fields (e.g., HCI and Biomedical), makes it challenging to find comprehensive and up-to-date literature. We also identified challenges in accessing domain-specific knowledge from experts. When seeking advice from peers, professors, or domain experts without already having an understanding of the domain, the guidance researchers receive is typically limited to being directed to research papers or resources, rather than receiving detailed explanations (“… it can be kind of hard to reach out and then try to explain what I want … some of the concepts require a lot of background knowledge and previous experience that I feel like they can’t really communicate to me just through an email, or like a 5 minutes talk … ,” P12). Participants also highlighted difficulties in finding the right person to consult, and there is often a gap in understanding between them and the experts they approach, as noted by P7 “… a gap of understanding between the two (researchers) … professors have different kind of understanding and expectation about our capabilities, they might be thinking that I might be aware of a statistical concept … but I do not know how to do it, and they’re like, but I thought you knew it already … .” The participants were also asked about their past experience using AI tools during the ideation stage when conducting interdisciplinary research. P10 mentioned that they would use ChatGPT for explaining unfamiliar concepts and terminology, while noting its tendency of being generic “… but ChatGPT sometimes can be confused … it sometimes would start repeating [the] same concepts … the responses are generic and not specific …” Other participants also shared similar remarks, including P12 “… hard to prompt it well enough to give more than generic responses, or sometimes it just gets too complicated.”
@@
  • [D2] Gathering feedback from the perspectives of domain experts: As highlighted by the formative study interviewees, in real-life research settings, getting feedback and sharing knowledge from domain experts is crucial for advancing novel ideas, but the process is often hindered by the lack of access and opportunities for communication to obtain timely and constructive input. Past theories (Mayo et al., 2017) have also suggested the significance of having access to the knowledge of a diverse set of expertise in the context of collaborative ideation. The system should have mechanisms for actively seeking and incorporating feedback from domain experts with diverse background knowledge in a timely manner.